OpenStudy (anonymous):

(a) Let a denote the first term and r the common ratio of an exponential sequence. (i) show that the general rule for finding the sum, Sn' of the first n term of the sequence is give by (a(1-r^n))/(1-r),r≠1. (ii) Hence deduce the sum of infinity of an exponential sequence with |r|<1.

OpenStudy (anonymous):

Since \[ |r|<1 \]We have \[ n\to \infty \implies r^n\to0 \]

OpenStudy (anonymous):

is that all

OpenStudy (anonymous):

You still must do part (i)

OpenStudy (anonymous):

could you please solve

OpenStudy (anonymous):

\[ (1-r)\sum_{k=0}^{n-1} r^k = \sum_{k=0}^{n-1} r^k -r^{k+1} = 1-r^n \]So \[ \sum_{k=0}^{n-1} r^k=\frac{1-r^n}{1-r} \]

OpenStudy (anonymous):

You make it a telescoping sum

OpenStudy (anonymous):

what do you imply by telescoping sum

OpenStudy (anonymous):

For the terms \(2,3,\dots , n-1\) they cancel each other out becaues you have \(r^0-r^1+r^1-r^2+r^2-\dots\)

OpenStudy (anonymous):

What next?